Georgia, Inc. has collected the following data on one of its products. The direct materials quantity variance IS. Direct materials standard (3 lbs $1/1b) Total direct materials cost variance-unfavorable $24,250 $3 per finished unit Actual direct materials used 108,000 lbs. 27,000 units Actual finished units produced Multiple Choice $24,250 unfavorable. $27,000 favorable. $2,750 favorable. $24,250 favorable. $27,000 unfavorable.
A company has an overhead application rate of 128% of direct labor costs. How much overhead would be allocated to a job if it required total labor costing $23,000? Multiple Choice $14,720 $17,969 $29,440 $294,400 $23,000
During March, a firm expects its total sales to be $156,000, its total variable costs to be $94,600, and its total fixed costs to be $24,600. The contribution margin for March is: Multiple Choice $36,800 $24,600. $119,200 $61,400. $94,600
Answer #1

Q1 As per data provided, 24500 unfavourable is the direct material variance

Q2:As per data provided, overhead = 128%*direct labour = 1.28*23000=$29440

Q3:Contribution margin = total sales - variablel cost = 156000-94600=$61400